# appendixReferences

### v1

@@ 3.2 Variational Inference
To reduce the model parameters, we share the encoder between the generative and inference processes. Therefore, $\bm{\tilde{\mu}}_{s}^{1:C}$ and $\bm{\tilde{\mu}}_{t}^{1:C}$ are computed via the same process described by Equation 1. In rule selection, RAISE reformulates the variational posterior of the indicator as $q(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})\propto p(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})p(r^{c}|\bm{z}_{S}^{c})$. We first predict the prior categorical probability $\bm{\pi}_{1:K}^{c}$ of $p(r^{c}|\bm{z}_{S}^{c})$ from the context $\bm{z}_{S}^{c}$. Then we predict the target concepts by executing each atomic rule on $\bm{z}_{S}^{c}$ and compute the likelihood $p(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})$ of each atomic rule. $q(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})$ is estimated by considering both the prior categorical probability and likelihood comprehensively, which reduces the risk of model collapse (e.g., the model always selects one rule). We provide the derivation of $q(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})$ in Appendix A.1. Letting $\Psi=\{\theta,\phi,\psi\}$, we factorize the variational distribution as
$q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})=\prod_{c=1}^{C}q_{\phi,\psi}(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})\prod_{s\in S}q_{\theta}(\bm{z}_{s}^{c}|\bm{x}_{s})\prod_{t\in T}q_{\theta}(\bm{z}_{t}^{c}|\bm{x}_{t}).$ (8)
@@ 3.3 Parameter Learning
We update the parameters of RAISE by maximizing the evidence lower bound (ELBO) to approximately optimize the log-likelihood $\log p(\bm{x}_{T}|\bm{x}_{S})$ Kingma & Welling (2013). With the generative process $p_{\Theta}$ and the variational distribution $q_{\Psi}$ defined in Equations 6 and 8, the ELBO is
@@
$=\underbrace{\sum_{t\in T}\mathbb{E}_{q}\big{[}\log p(\bm{x}_{t}|\bm{z}_{t})\big{]}}_{\text{$\mathcal{L}_{\text{rec}}$}}-\underbrace{\sum_{c=1}^{C}\mathbb{E}_{q}\left[\log\frac{\prod_{t\in T}q(\bm{z}_{t}^{c}|\bm{x}_{t})}{p(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})}\right]}_{\text{$\mathcal{R}_{\text{pred}}$}}-\underbrace{\sum_{c=1}^{C}\mathbb{E}_{q}\left[\log\frac{q(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})}{p(r^{c}|\bm{z}_{S}^{c})}\right]}_{\text{$\mathcal{R}_{\text{rule}}$}}$
where $q$ denotes $q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})$, and the parameter symbols are omitted for convenience. $\mathcal{L}_{\text{rec}}$ is reconstruction loss that measures the quality of the reconstruction images. The concept regularizer $\mathcal{R}_{\text{pred}}$ estimates the distance between the predicted target concepts and those directly encoded from target images. Minimizing $\mathcal{R}_{\text{pred}}$ will promote RAISE to generate correct predictions of latent concepts. The rule regularizer $\mathcal{R}_{\text{rule}}$ makes RAISE select the same rules when given different context images of a matrix. The variational posterior $q_{\phi,\psi}(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})$ conditioned on the entire matrix and the prior $p_{\phi}(r^{c}|\bm{z}_{S}^{c})$ conditioned on the context images are expected to have similar probabilities. The detailed derivation of the ELBO is provided in Appendix A.2.
As the rule abstraction and selection are based on concept learning, RAISE introduces auxiliary rule annotations to help stabilize the learning of latent concepts. A rule annotation $\bm{v}=\{v_{a}\}_{a=1}^{A}$ contains indicator variables of $A$ ground truth attributes to illustrate the rule on each attribute. For example, If there are three ground truth attributes, the rule annotation $\bm{v}=[1,0,2]$ where each element indicates the rule types on one attribute. In training, RAISE will not leverage the meta information of attributes. RAISE is informed of the rule types on three attributes, but does not know what the attributes represent. Therefore, we need to determine the correspondence between latent concepts and attributes, so that each ground truth indicator can guide the learning of related latent concepts. RAISE introduces a $A\times C$ binary matrix $\bm{M}$ where $\bm{M}_{a,c}=1$ means that the $a$th attribute is related to the $c$th concept. An auxiliary loss is used to measure the accuracy of rule selection:
@@
$\mathcal{L}_{\text{sup}}$ is the log-likelihood of the category distributions considering the attribute-concept correspondence $\bm{M}$. The binary matrix $\bm{M}$ is determined by solving
$\begin{gathered}\operatorname*{arg\,max}_{\bm{M}}\mathcal{L}_{\text{sup}}\quad\text{s.t.}\quad\begin{cases}\sum_{c=1}^{C}\bm{M}_{a,c}=1,&a=1,...,A,\\ \sum_{a=1}^{A}\bm{M}_{a,c}=0\text{ or }1,&c=1,...,C,\\ \bm{M}_{a,c}=0\text{ or }1,&a=1,...,A,\quad c=1,...,C.\end{cases}\end{gathered}$ (11)
The assignment problem can be solved by the modified Jonker-Volgenant algorithm Crouse (2016). Equation 11 handles the redundancy of concepts by allowing some latent concepts not to encode any attribute. Including the auxiliary loss, the training objective is
@@
where $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are hyperparameters to control the importance of subterms. The training of RAISE supports semi-supervised settings. For samples that do not provide rule annotations, RAISE can set $\beta_{3}=0$ and update parameters via the unsupervised part $\mathcal{L}_{\text{rec}}-\beta_{1}\mathcal{R}_{\text{pred}}-\beta_{2}\mathcal{R}_{\text{rule}}$.
@@ 4 Experiments
Datasets. The models in the experiments are evaluated on the RAVEN Zhang et al. (2019a) and I-RAVEN Hu et al. (2021) datasets having seven image configurations (e.g., scenes with centric objects or object grids) and four basic rules. I-RAVEN follows the same configurations as RAVEN and reduces the bias of candidate sets to resist the shortcut learning of models Hu et al. (2021). See Appendix B for detailed descriptions of datasets.
Compared Models. In the bottom-right answer selection task, we compare RAISE with the powerful generative solvers ALANS Zhang et al. (2021b), PrAE Zhang et al. (2021a), and the model proposed by Niv et al. (called GCA for convenience) Pekar et al. (2020). RAISE selects the candidate closest to the predicted result in the latent space as the answer. We apply three strategies of answer selection in GCA: selecting the candidate having the smallest pixel difference to the prediction (GCA-I), having the smallest difference in the representation space (GCA-R), and having the highest panel score (GCA-C). Since these generative solvers cannot generate non-bottom-right answers, we take Transformer Vaswani et al. (2017), ANP Kim et al. (2019), LGPP Shi et al. (2021), and CLAP Shi et al. (2023) as baseline models to evaluate the ability to generate answers at arbitrary positions. We give more detail about the related models in Appendix C.
Training and Evaluation Settings. For non-grid layouts, RAISE is trained under semi-supervised settings by using 5% rule annotations. RAISE leverages 20% rule annotations on O-IG and full rule annotations on 2$\times$2Grid and 3$\times$3Grid. The powerful generative solvers use full rule annotations and are trained and tested on each configuration respectively. We compare RAISE with them to illustrate the acquired bottom-right answer selection ability of RAISE under semi-supervised settings. The baselines can generate answers at arbitrary positions but cannot leverage rule annotations since they do not explicitly model the category of rules. We compare RAISE with the baselines to illustrate the benefit of learning latent concepts and atomic rules for generative abstract reasoning. Since the training of RAISE and the baselines do not require the candidate sets, and RAVEN/I-RAVEN only differ in the distribution of candidates, we train RAISE and the baselines on RAVEN and test them on both RAVEN/I-RAVEN directly. See Appendix C for detailed training and evaluation settings.
@@ 4.1 Bottom-Right Answer Selection
Table 1: The accuracy (%) of selecting bottom-right answers on RAVEN/I-RAVEN. For models with randomness, the table displays the average results of ten trials.
@@
A classic RPM test requires participants to select an answer from the candidate set to fill in the missing image at the bottom right. The results in Table 1 illustrate RAISE’s outstanding generative abstract reasoning ability. By comparing the difference between predictions and candidates, RAISE outperforms the compared generative solvers in most configurations of RAVEN and I-RAVEN, even if the distractors in candidate sets are not used in training. All the powerful generative solvers take full rule annotations for training, while RAISE in non-grid configurations only requires a small amount of rule annotations (5% samples) to achieve high selection accuracy. RAISE attains the highest selection accuracy compared to the baselines which can generate answers at arbitrary positions. By comparing the results on RAVEN and I-RAVEN, we find that the generative solvers are more likely to have accuracy improvement on I-RAVEN. I-RAVEN generates distractors less similar to correct answers than RAVEN to avoid significant biases in candidate sets, making answer selection easier. For grid-shaped configurations, we found that the noise in datasets will significantly influence the model performance. By removing the noise in object attributes, RAISE achieves high selection accuracy on three grid-shaped configurations using only 20% rule annotations. See Appendix D.1 for the detailed descriptions and experimental results.
4.2 Answer Selection at Arbitrary Position
@@
Figure 3: Answer generation at arbitrary positions. We give the predictions on RAVEN with red boxes to illustrate the ability of arbitrary-position answer generation. Some predictions of RAISE differ from the original sample, but they still follow the correct rules due to the existence of noise.
The above generative solvers can hardly generate answers at non-bottom-right positions. In this experiment, we probe the ability of RAISE and baselines to generate answers at arbitrary positions, which is acquired by predicting random target images in RPMs. We first generate additional candidate sets in the experiment because RAVEN and I-RAVEN do not provide candidate sets for non-bottom-right images. To this end, we sample a batch of RPMs from the dataset and split the RPMs into target and context images in the same way. For each matrix, we use the target images of other $N_{c}$ samples in the batch as distractors to generate a candidate set with $N_{c}+1$ entries. This strategy can adapt to the missing images at arbitrary and even multiple positions, and we can easily control the difficulty of answer selection through the number of distractors
@@ 4.3 Latent Concepts
Latent concepts bridge atomic rules and high-dimensional observations. Figure 4a visualizes the latent concepts learned from Center and O-IC by traversing concept representations of an image in the latent space. If the concepts are well decomposed, decoding the interpolated concept representations will change one attribute of the original image. Besides observing visualization results, we can find the correspondence between concepts and attributes with the aid of the binary matrix $\bm{M}$. As shown in Figure 4a, RAISE can automatically set some redundant concepts when there are more concepts than attributes. (e.g., the 1st, 3rd, 5th, 6th, and 8th concepts of the configuration Center). The visualization results illustrate the concept learning ability of RAISE, which is the foundation of abstracting and selecting global atomic rules shared among RPMs.
@@ 4.4 Odd-One-Out in RPM
In odd-one-out tests, RAISE attempts to pick up the image that breaks the underlying rule of a panel. To generate RPM-based odd-one-out problems, we replace the bottom-right image of an RPM with a random distractor in the candidate set. Taking Figure 4b as an example, by replacing the bottom-right image, we change the object color from white to black. RAISE takes each image in the RPM as the target, gets the prediction results, and computes the distance in concepts between the predictions and original images. The right panel of Figure 4b shows the concept distances of each image, and we find that the 7th concept of the bottom-right image deviates the most. According to Figure 4a, the 5th concept on Center represents the attribute Color (i.e., the type of inner object), which is indeed the attribute modified when constructing the test. The last row has relatively higher concept distances since the incorrect image tends to influence the accuracy of answer generation at the most related positions. However, because the concepts and concept-specific rules are designed as independent in RAISE, the high concept distances only appear in the 7th concept. By solving RPM-based odd-one-out problems in this experiment, we explain how concept-level predictions improve the interpretability of answer selection. Though RAISE is tasked with generating answers, it can handle answer-selection problems by excluding candidates that violate the underlying rules.

### v3

@@ 3.2 Variational Inference
Since RAISE shares the encoder between the generative and inference processes to reduce the model parameters, we compute context latent concepts $\bm{\tilde{\mu}}_{s}^{1:C}$ and target latent concepts $\bm{\tilde{\mu}}_{t}^{1:C}$ via the same process described in Equation 1. In the inference process, RAISE reformulates the variational distribution of the categorical indicator $r^{c}$ as $q(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})\propto p(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})p(r^{c}|\bm{z}_{S}^{c})$. That is, RAISE predicts the prior probabilities $\bm{\pi}_{1:K}^{c}$ of $p(r^{c}|\bm{z}_{S}^{c})$ from the context latent concepts $\bm{z}_{S}^{c}$ and compute the likelihood $p(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})$ by executing the atomic rule $r^{c}$ ($r^{c}=1,\cdots,K$) on $\bm{z}_{S}^{c}$. In this way, we can estimate the variational distribution $q(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})$ by considering both the prior probabilities and the likelihoods of $K$ atomic rules, which reduces the risk of model collapse (e.g., always selecting one atomic rule from $\psi$). We provide more details of $q(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})$ in Appendix A.1. Letting $\Psi=\{\theta,\phi,\psi\}$, we factorize the variational distribution as
$q_{\Psi}(\bm{h}|\bm{x}_{T},\bm{x}_{S})=\prod_{c=1}^{C}\left(q_{\phi,\psi}(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})\prod_{s\in S}q_{\theta}(\bm{z}_{s}^{c}|\bm{x}_{s})\prod_{t\in T}q_{\theta}(\bm{z}_{t}^{c}|\bm{x}_{t})\right).$ (8)
@@ 3.3 Parameter Learning
We update the parameters of RAISE by maximizing the evidence lower bound (ELBO) of the log-likelihood $\log p(\bm{x}_{T}|\bm{x}_{S})$ Kingma & Welling (2013). With the generative process $p_{\Theta}$ and the variational distribution $q_{\Psi}$ defined in Equations 6 and 8, the ELBO is ($q$ denotes the variational distribution, and we omit the parameter symbols $\Theta$ and $\Psi$ for convenience)
@@
$=\underbrace{\sum_{t\in T}\mathbb{E}_{q}\big{[}\log p(\bm{x}_{t}|\bm{z}_{t})\big{]}}_{\text{$\mathcal{L}_{\text{rec}}$}}-\underbrace{\sum_{c=1}^{C}\mathbb{E}_{q}\left[\log\frac{q(\bm{z}_{T}^{c}|\bm{x}_{T})}{p(\bm{z}_{T}^{c}|r^{c},\bm{z}_{S}^{c})}\right]}_{\text{$\mathcal{R}_{\text{pred}}$}}-\underbrace{\sum_{c=1}^{C}\mathbb{E}_{q}\left[\log\frac{q(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})}{p(r^{c}|\bm{z}_{S}^{c})}\right]}_{\text{$\mathcal{R}_{\text{rule}}$}}$
The reconstruction loss $\mathcal{L}_{\text{rec}}$ measures the quality of the reconstruction images. The concept regularizer $\mathcal{R}_{\text{pred}}$ estimates the distance between the predicted target concepts and the concepts directly encoded from target images. Minimizing $\mathcal{R}_{\text{pred}}$ will promote RAISE to generate correct predictions in the space of latent concepts. The rule regularizer $\mathcal{R}_{\text{rule}}$ expects RAISE to select the same rules when given different sets of images in an RPM. The variational posterior $q(r^{c}|\bm{z}_{S}^{c},\bm{z}_{T}^{c})$ conditioned on the entire matrix and the prior $p(r^{c}|\bm{z}_{S}^{c})$ conditioned on the context images are expected to have similar probabilities. The detailed derivation of the ELBO is provided in Appendix A.2.
The abstraction and selection of atomic rules rely on the acquired latent concepts. Therefore, RAISE introduces auxiliary rule annotations to improve the quality of latent concepts and stabilize the learning process. We denote rule annotations as $\bm{v}=\{v_{a}\}_{a=1}^{A}$ where $A$ is the number of ground truth attributes and $v_{a}$ indicates the type of rules on the $a$-th attribute. For example, $\bm{v}=[2,1,3]$ means that the attributes follow the second, first, and third rules respectively. RAISE does not leverage the meta-information of attributes in training since the rule annotations only inform the type of rule on each attribute. The meaning of attributes is automatically learned by RAISE for accurate rule abstraction and selection. One key to guiding concept learning with rule annotations is determining the correspondence between latent concepts and attributes. RAISE introduces a $A\times C$ binary matrix $\bm{M}$ where $\bm{M}_{a,c}=1$ indicates that the $a$-th attribute is encoded in the $c$-th latent concept. Therefore, the rule predicted on the $c$-th latent concept is supervised by the rule annotation $v_{a}$, and the auxiliary loss measures distances between the predicted and ground truth types of rules:
@@
The auxiliary loss $\mathcal{L}_{\text{sup}}$ is the log-likelihood of the categorical distributions considering the attribute-concept correspondence $\bm{M}$. The binary matrix $\bm{M}$ is derived by solving the following assignment problem on a batch of RPM samples:
$\begin{gathered}\operatorname*{arg\,max}_{\bm{M}}\text{ }\mathcal{L}_{\text{sup}}\quad\text{s.t.}\quad\begin{cases}\sum_{c=1}^{C}\bm{M}_{a,c}=1,&a=1,...,A,\\ \sum_{a=1}^{A}\bm{M}_{a,c}=0\text{ or }1,&c=1,...,C,\\ \bm{M}_{a,c}=0\text{ or }1,&a=1,...,A,\quad c=1,...,C.\end{cases}\end{gathered}$ (11)
Equation 11 allows the existence of redundant latent concepts, which can be solved using the modified Jonker-Volgenant algorithm Crouse (2016). In this case, the training objective becomes
@@
where $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are hyperparameters. RAISE also supports semi-supervised training settings. For samples that do not provide rule annotations, RAISE can set $\beta_{3}=0$ and update parameters via the unsupervised part $\mathcal{L}_{\text{rec}}-\beta_{1}\mathcal{R}_{\text{pred}}-\beta_{2}\mathcal{R}_{\text{rule}}$.
@@ 4 Experiments
Datasets. The models in the experiments are evaluated on the RAVEN Zhang et al. (2019a) and I-RAVEN Hu et al. (2021) datasets having seven image configurations (e.g., scenes with centric objects or object grids) and four basic rules. I-RAVEN follows the same configurations as RAVEN and reduces the bias of candidate sets to resist the shortcut learning of models Hu et al. (2021). See Appendix B for details of datasets.
Compared Models. In the task of bottom-right answer selection, we compare RAISE with the powerful generative solvers ALANS Zhang et al. (2021b), PrAE Zhang et al. (2021a), and the model proposed by Niv et al. (called GCA for convenience) Pekar et al. (2020). RAISE selects the candidate closest to the predicted result in the latent space as the answer. We apply three strategies of answer selection in GCA: selecting the candidate having the smallest pixel difference to the prediction (GCA-I), having the smallest difference in the representation space (GCA-R), and having the highest panel score (GCA-C). Since these generative solvers cannot generate non-bottom-right answers, we take Transformer Vaswani et al. (2017), ANP Kim et al. (2019), LGPP Shi et al. (2021), and CLAP Shi et al. (2023) as baseline models to evaluate the ability to generate answers at arbitrary positions. We provide more details in Appendix C.
Training and Evaluation Settings. For non-grid layouts, RAISE is trained under semi-supervised settings by using 5% rule annotations. RAISE leverages 20% rule annotations on O-IG and full rule annotations on 2$\times$2Grid and 3$\times$3Grid. The powerful generative solvers use full rule annotations and are trained and tested on each configuration respectively. We compare RAISE with them to illustrate the acquired bottom-right answer selection ability of RAISE under semi-supervised settings. The baselines can generate answers at arbitrary positions but cannot leverage rule annotations since they do not explicitly model the category of rules. We compare RAISE with the baselines to illustrate the benefit of learning latent concepts and atomic rules for generative abstract reasoning. Since the training of RAISE and the baselines do not require the candidate sets, and RAVEN/I-RAVEN only differ in the distribution of candidates, we train RAISE and the baselines on RAVEN and test them on RAVEN/I-RAVEN directly. See Appendix C for detailed training and evaluation settings.
@@ 4.1 Bottom-Right Answer Selection
Table 1: The accuracy (%) of selecting bottom-right answers on different configurations (i.e., Center, L-R, etc) of RAVEN/I-RAVEN. The table displays the average results of ten trials.
@@
This experiment conducts classical RPM tests that require models to find the missing bottom-right images in eight candidates. Table 1 illustrates RAISE’s outstanding generative abstract reasoning ability on RAVEN/I-RAVEN. By comparing the difference between predictions and candidates, RAISE outperforms the compared generative solvers in most configurations of RAVEN/I-RAVEN, even if the distractors in candidate sets are not used in training. All the powerful generative solvers take full rule annotations for training, while RAISE in non-grid configurations only requires a small amount of rule annotations (5% samples) to achieve high selection accuracy. RAISE attains the highest selection accuracy compared to the baselines which can generate answers at arbitrary positions. By comparing the results on RAVEN/I-RAVEN, we find that generative solvers are more likely to have accuracy improvement on I-RAVEN, because I-RAVEN generates distractors that are less similar to correct answers to avoid significant biases in candidate sets. For grid-shaped configurations, we found that the noise in datasets will significantly influence the model performance. By removing the noise in object attributes, RAISE achieves high selection accuracy on three grid-shaped configurations using only 20% rule annotations. See Appendix D.1 for the detailed experimental results.
4.2 Answer Selection at Arbitrary Positions
@@
Figure 3: Answer generation at arbitrary positions. The prediction results on RAVEN are highlighted (red box) to illustrate the arbitrary-position generation ability. Due to the existence of noise, some predictions may differ from the original sample, but they still follow the correct rules.
The above generative solvers can hardly generate answers at non-bottom-right positions. In this experiment, we probe the ability of RAISE and baselines to generate answers at arbitrary positions. We first generate additional candidate sets in the experiment because RAVEN and I-RAVEN do not provide candidate sets for non-bottom-right images. To this end, we sample a batch of RPMs from the dataset and split the RPMs into target and context images in the same way. For each matrix, we use the target images of other $N_{c}$ samples in the batch as distractors to generate a candidate set with $N_{c}+1$ entries. This strategy can adapt to the missing images at arbitrary and even multiple positions, and we can easily control the difficulty of answer selection through the number of distractors
@@ 4.3 Latent Concepts
Latent concepts bridge atomic rules and high-dimensional observations. Figure 4a visualizes the latent concepts learned from Center and O-IC by traversing concept representations of an image in the latent space. If the concepts are well decomposed, decoding the interpolated concept representations will change one attribute of the original image. Besides observing visualization results, we can find the correspondence between concepts and attributes with the aid of the binary matrix $\bm{M}$. As shown in Figure 4a, RAISE can automatically set some redundant concepts when there are more concepts than attributes. (e.g., the first concept of Center). The visualization results illustrate the concept learning ability of RAISE, which is the foundation of abstracting and selecting atomic rules shared among RPMs.
@@ 4.4 Odd-One-Out in RPM
In odd-one-out tests, RAISE attempts to find the rule-breaking image in a panel. To generate RPM-based odd-one-out problems, we replace the bottom-right image of an RPM with a random distractor in the candidate set. Taking Figure 4b as an example, we change the object color from white to black by replacing the bottom-right image. RAISE takes each image in an RPM as the target, gets the prediction results, and computes the prediction error on latent concepts. The right panel of Figure 4b shows the concept-level prediction errors, and we find that the 7th concept of the bottom-right image deviates the most. According to Figure 4a, the 7th concept on Center represents the attribute Color, which is indeed the attribute modified when constructing the test. The last row has relatively higher concept distances since the incorrect image tends to influence the accuracy of answer generation at the most related positions. Because of the independent latent concepts and concept-specific reasoning processes of RAISE, the high concept distances only appear in the 7th concept. By solving RPM-based odd-one-out problems, we explain how concept-level predictions improve the interpretability of answer selection. Although RAISE is tasked with generating answers, it can handle answer-selection problems by excluding candidates violating the underlying rules.